Find the capacitive reactance of a 20 microF capacitor connected to a 500 Hz source.

  1. Find the capacitive reactance of a 20 microF capacitor connected to a 500 Hz source.
Answer #1

Capacitive reactance is given by,

Xc = 1/(w*C)

here, C = capacitance = 20 \mu F = 20*10^-6 F

w = angular frequency = 2*pi*f

f = frequency = 500 Hz

then,

Xc = 1/[2*pi*500*(20*10^-6)]

Xc = 15.9 ohm

"Let me know if you have any query."